# Quantum criticality and universality in stationary state of long-range Kitaev model

### v1

@@ IV.1 Signature of criticality in stationary state
As we look for the signature of quantum criticality in the long-time stationary state for different $\alpha$ values, we consider two types of quench protocols which differ in the initial states. In the first protocol, we consider initial state as the ground state of LRK Hamiltonian corresponding to the parameter values $\mu_{i}=1$ and $\Delta_{i}=-1$ and will be referred henceforth as critical state. The second protocol correspond to the initial state as ground state of LRK Hamiltonian corresponding to parameters $\mu_{i}=1.5$ and $\Delta_{i}=-1$ which will be referred henceforth as non-critical state. The final Hamiltonian parameters, $\mu_{f}$ and $\Delta_{f}$, are varied to include both the critical and non-critical regimes. Note that it has already been established that mutual information of ground state shows a peak at the critical point namely $\mu=\pm 1$ for large $\alpha(=10)$ while only at $\mu=1$ for $\alpha\leq 1$ [71]. Mutual information, $I_{A_{1}:A_{2}}$ calculated for stationary state for both quench protocols i.e. $(\mu_{i}=1,\Delta_{i}=-1)\to(\mu_{f}=\mu,\Delta_{f}=1)$ and $(\mu_{i}=1.5,\Delta_{i}=-1)\to(\mu_{f}=\mu,\Delta_{f}=1)$ are plotted in Fig.1 (row 1). For final quench parameters corresponding to critical values, a peak in $I_{A_{1}:A_{2}}$ for critical initial state is clearly visible indicating the signature of the criticality in stationary state for all values of $\alpha$. In the other quench protocol with non-critical initial state, $I_{A_{1}:A_{2}}$ does show a non-analytic behaviour in the form of a sharp dip for $\mu_{f}=1$ for all values of $\alpha$. To confirm that this behaviour is indeed due to quantum nature of correlations, we plot log-negativity $\xi_{A_{1}:A_{2}}$ in Eq. 11 in Fig. 1 (row 2) for the same quench protocols. The similar peak and dip structure is unmistakably seen in behaviour of $\xi_{A_{1}:A_{2}}$ whenever the final Hamiltonian is critical. Such behaviour of $I_{A_{1}:A_{2}}$ for nearest neighbour free-fermionic model has earlier been explained in terms of soft mode for which the energy vanishes near criticality. For a critical-to-critical quench, the soft mode does not get excited resulting in a peak structure in $I_{A_{1}:A_{2}}$ while it heats up to infinite temperature for a noncritical-to-critical quench displaying a dip in $I_{A_{1}:A_{2}}$. This further manifests as an algebraic decay of fermionic correlation for critical-to-critical quench, akin to ground-state behavior, while for the non-critical-to-critical quench, it decays exponentially indicating a non-critical behavior. However, such an explanation falls short for LRK model for $\alpha\leq 1$, as correlation decay pattern in the ground state is always algebraic.
@@
For this non-analytic behaviour of entanglement measure at $\mu_{f}=1$, we offer an alternate explanation in terms of the mode occupation probability, $n_{k}$, for soft mode ($k_{c}=\pi$) where the energy vanishes near criticality. Let’s recall that $n_{k_{c}}=1/2$ would imply contribution from multiple modes and therefore leads to higher entanglement entropy for union of two subsystems, $S_{vN}^{A_{1}\cup A_{2}}$. As in $I_{A_{1}:A_{2}}$ in Eq. 9, $S_{vN}^{A_{1}\cup A_{2}}$ is being subtracted, this leads to decrease in $I_{A_{1}:A_{2}}$. In contrast, $n_{k_{c}}=1,0$ will lead to reduction of $S_{vN}^{A_{1}\cup A_{2}}$ and consequently leads to increment of the mutual information.
@@
The occupation probability, $n_{k}=\langle\eta_{k}^{\dagger}\eta_{k}\rangle$ of the Bogoliubov modes, $\eta_{k}$, in the stationary state of LRK model is obtained as
@@
where $\lambda^{i}_{\alpha}\left(k\right)(\lambda^{f}_{\alpha}\left(k\right))$ denotes the energy corresponding to the pre(post)-quench Hamiltonian. From Eq. 23, it can be shown that if we start with the non-critical initial state, i.e., $\mu_{i}=1.5$, in the limit $k\to\pi$, the occupation probability. $n_{k}$ is given by
@@
as illustrated in Fig.1(row 3). This explains the dips observed in $I_{A_{1},A_{2}}$ in the long-time stationary state for a non-critical-to-critical quench. In contrast, for the quench protocol involving critical state, $n_{k}$ near the soft mode is
@@
This explains the peak observed in $I_{A_{1},A_{2}}$ for the critical-to-critical quench protocol.
Figure 2: Phase plot of effective central charge $c_{\rm eff}^{I}$ extracted from mutual information in the post-quench $\Delta-\mu$ plane for both critical (first column) and non-critical (second column) initial state for $\alpha=0$ (first row), 1 (second row) and 2 (third row). The parameter $\Delta_{i}$ is fixed at -1 for all the plots.
@@ IV.2 Signature of universality in stationary state
After establishing the non-analytic behaviour of mutual information (and log-negativity) for post-quench stationary state at critical point of post-quench Hamiltonian, a natural question arises whether information of effective central charge, $c_{\rm eff}^{I}$ ($c_{\rm eff}^{N}$) calculated from scaling of $I_{A_{1}:A_{2}}$ ($\xi_{A_{1}:A_{2}}$) depends on the quench protocol? Whether this effective central charge is universal i.e.independent of quench protocols in $\Delta$, if yes whether the central charge of ground state is conserved? To address these questions, we analyze both $c_{\rm eff}^{I}$ and $c_{\rm eff}^{N}$ for both the quench protocols discussed in sub-section IV.1. Here, it is useful to recall that central charge provides a powerful tool to understand the underlying CFT that describes the particular universality class [33]. For instance, the Ising universality class and the Luttinger liquid universality class are described by completely different CFT and characterized by distinct central charges. The central charge gives a measure of the number of degrees of freedom or the “size” of the symmetry in the theory [90; 91].
Fig. 2 displays $c^{I}_{\rm eff}$ as a function of the post-quench parameters $\mu_{f}$ and $\Delta_{f}$. The columns of Fig. 2 represents $c^{I}_{\rm eff}$ for quenches starting with two distinct initial states, while rows correspond to varying range of $\alpha$. Let’s focus on the first column i.e. quench from critical state for different values of $\alpha$. For both $\alpha=0$ and $\alpha=2$, $c_{\rm eff}^{I}$ is independent of $\Delta$-quenches and retain the same value as the ground state for a critical-to-critical quench protocol. For other values of $\mu_{f}$, $c_{\rm eff}^{I}$ calculated from stationary state takes the value of non-critical ground state, i.e. $0.5$ for $\alpha=0$ and 0 for $\alpha=2$  [^1] Note that $\Delta=0$ completely changes the Hamiltonian to a nearest neighbour model with no pairing term with effective central charge 0 for $\mu\neq 1$.. For $\alpha=1$, the effective central charge for ground state at critical point depends on the value of $\Delta$ and this character is retained in the post-quench stationary state as well. Therefore, the universality can only be discussed in the case of $\alpha=0$ and $\alpha=2$ from a critical-to-critical quench. It is clearly seen that $c_{\rm eff}^{I}$ calculated from post-quench stationary state conserves the central charge (see Table-1). This implies that universality class for the critical ground state can also be inferred from stationary state.
Table 2: Summary of effective central charge values calculated from post-quench stationary state for critical-to-critical, noncritical-to-critical and critical-to-noncritical quenches of LRK model for different values of $\alpha$.
@@
Now for the second column of Fig. 2, universal characteristic of the $c_{\rm eff}^{I}$ of stationary state is reflected through independence of $\Delta$-quenches for $\alpha=0$ and $\alpha=2$ while $\alpha=1$ remains the $\Delta$-dependent. Note that even though entanglement measures showed a non-analytic behaviour at $\mu_{f}=1$ for non-critical initial state, the effective central charge does not show any such transition for either $\alpha=0$ or 2. Therefore, despite the non-analytic behaviour of entanglement measures at $\mu_{f}=1$, we do not attribute as signature of criticality or an universal description in terms of underlying CFT.
Therefore, to summarize, signature of universality in terms of effective central charge is seen in post-quench stationary state for a critical-to-critical quench. The effective central charges values in this case is summarized in tabular form in Table 2 which are in agreement with the values in Table 1 calculated for ground state.
Now, to investigate whether the long-range correlations in the stationary state are genuinely quantum in nature, we have evaluated the effective central charge, $c_{\rm eff}^{N}$, extracted from the finite-size scaling of the log-negativity for the stationary state. While the details are in Appendix A, the findings of effective central charge behaviour and therefore universality remains same as calculated using mutual information.
@@ V Summary
In summary, we studied the quench dynamics of the long-range Kitaev chain to investigate the possible signatures of quantum criticality in the long-time stationary state for different values of the exponent $\alpha$ characterizing the range of $p$-wave pairing interaction. To probe the criticality, we consider the quantum information-based measures such as mutual information and log-negativity. Our results show that a peak emerges in both mutual information and log-negativity for a critical-to-critical quench for all values of $\alpha$ studied in this work, exhibiting behavior similar to that of the critical ground state. This clearly indicates the presence of criticality in the stationary state. Importantly, the appearance of this peak can not be attributed to the change in decay pattern of fermionic correlation from exponential to algebraic at the critical point, as the fermionic correlators decay algebraically even at non-critical points for $\alpha\leq 1$. We argue that the peak arises as a result of the change in the occupation probability of the Bogoliubov fermions in the soft mode at the critical point.
We further show that for $\alpha=2$, long-range correlations develop only for the critical-to-critical quench. This result is obtained by analyzing the effective central charges, which are extracted from the scaling analysis of both mutual information and log-negativity. The effective central charge of the corresponding stationary state matches with the same for the critical ground state, regardless of the quench in the strength of the $p$-wave pairing interaction, $\Delta$. This clearly indicates that universality class for the critical ground state can also be inferred from stationary state. In contrast, for $\alpha=0$, the effective central charge is non-zero for all quench protocols, leading to long-range correlations, except when the final Hamiltonian parameter, $\Delta=0$. When $\Delta\neq 0$, the effective central charge is $1$ only for the critical-to-critical quench, and is $1/2$ for all other quench protocols, consistent with the ground state. This infers that the stationary state for critical to critical quench protocol belongs to the same universality class as the ground state.
For $\alpha=1$, the stationary state, similar to the ground state, cannot be described by any universality class, as the presence or absence of long-range correlations depends on the $\Delta$ quenches. However, long-range correlations do develop for the critical-to-critical quench protocol, regardless of the $\Delta$ quenches.
@@ VI Acknowledgment
Appendix A Overview of central charge for post-quench stationary state from log-negativity
Starting with the critical initial state, as shown in first row left column of Fig. 3, we find that for $\alpha=0$, $c_{\rm eff}^{N}=1/2$ for $\mu_{f}\neq 1$ and $\Delta_{f}\neq 0$, while $c_{\rm eff}^{N}=1$ for $\mu_{f}=1$ and $\Delta_{f}\neq 0$. In contrast, starting with a non-critical initial state, as shown in first row right column of Fig. 3, $c_{\rm eff}^{N}=1/2$ for all values of $\mu_{f}$ and $\Delta_{f}$, except for $\Delta_{f}=0$. This implies that for $\alpha=0$, $c_{\rm eff}^{N}=c_{\rm eff}^{I}$ for all quench protocols. For $\alpha=2$, similar to $c_{\rm eff}^{I}$, $c_{\rm eff}^{N}$ is non-zero and equal to $1/2$ only for a critical-to-critical quench protocol (with the exception of $\Delta_{f}=0$, where $c_{\rm eff}^{N}=0$). For all other quench protocols, $c_{\rm eff}^{N}=0$. This confirms that the long-range correlations associated with a non-zero effective central charge for both $\alpha=0$ and $\alpha=2$ in the stationary state are strictly quantum in nature. On the other hand, for $\alpha=1$, long-range correlations develop irrespective of $\Delta$ quenches in the critical-to-critical quench protocol (see second row first column of Fig. 3). Interestingly, for this quench protocol, $c_{\rm eff}^{N}$ is slightly larger than $c_{\rm eff}^{I}$. In contrast, for the critical-to-non-critical quench protocol, the presence of long-range correlations in the stationary state depends on the $\Delta$ quenches. Specifically, if $\Delta_{f}\Delta_{i}\gtrsim 0$, both $c_{\rm eff}^{I}$ and $c_{\rm eff}^{N}$ are non-zero, whereas both are zero if $\Delta\Delta_{0}\lesssim 0$. For all the quench protocols where $\Delta_{f}\Delta_{i}\gtrsim 0$, $c_{\rm eff}^{N}>c_{\rm eff}^{I}$.
Figure 3: Phase plot of the effective central charge from log-negativity $c_{\rm eff}^{N}$ for both critical (first column) and non-critical (second column) initial state for $\alpha=0$ (first row), 1 (second row) and 2 (third row). The quench parameters are same as for Fig. 2.

### v2

@@ IV.1 Signature of criticality in stationary state
As we look for the signature of quantum criticality in the long-time stationary state for different $\alpha$ values, we consider two types of quench protocols which differ in the initial states. In the first protocol, we consider the initial state as the ground state of the LRK Hamiltonian corresponding to the parameter values $\mu_{i}=1$ and $\Delta_{i}=-1$ and is henceforth referred to as the critical state. The second protocol corresponds to the initial state as the ground state of the LRK Hamiltonian corresponding to parameters $\mu_{i}=1.5$ and $\Delta_{i}=-1$, which is henceforth referred to as the noncritical state. The final Hamiltonian parameters, $\mu_{f}$ and $\Delta_{f}$, are varied to include both the critical and noncritical regimes. Note that it has already been established that mutual information of the ground state shows a peak at the critical point, namely, $\mu=\pm 1$ for large $\alpha(=10)$, while only at $\mu=1$ for $\alpha\leq 1$ [71]. Mutual information $I_{A_{1}:A_{2}}$ calculated for the stationary state for both quench protocols, i.e., $(\mu_{i}=1,\Delta_{i}=-1)\to(\mu_{f}=\mu,\Delta_{f}=1)$ and $(\mu_{i}=1.5,\Delta_{i}=-1)\to(\mu_{f}=\mu,\Delta_{f}=1)$, is plotted in Fig. 1 (row 1). For the final quench parameters corresponding to critical values, a peak in $I_{A_{1}:A_{2}}$ for the critical initial state is clearly visible, indicating the signature of the criticality in the stationary state for all values of $\alpha$. In the other quench protocol with the noncritical initial state, $I_{A_{1}:A_{2}}$ does show a nonanalytic behavior in the form of a sharp dip for $\mu_{f}=1$ for all values of $\alpha$. To confirm that this behavior is indeed due to the quantum nature of correlations, we plot log-negativity $\xi_{A_{1}:A_{2}}$ in Eq. 11 in Fig. 1 (row 2) for the same quench protocols. The similar peak and dip structure is unmistakably seen in the behavior of $\xi_{A_{1}:A_{2}}$ whenever the final Hamiltonian is critical. Such behavior of $I_{A_{1}:A_{2}}$ for the nearest-neighbor free-fermionic model has earlier been explained in terms of a soft mode for which the energy vanishes near criticality. For a critical-to-critical quench, the soft mode does not get excited resulting in a peak structure in $I_{A_{1}:A_{2}}$ while it heats up to infinite temperature for a noncritical-to-critical quench displaying a dip in $I_{A_{1}:A_{2}}$. This further manifests as an algebraic decay of fermionic correlation for critical-to-critical quench, akin to ground-state behavior, while for the noncritical-to-critical quench, it decays exponentially, indicating a noncritical behavior. However, such an explanation falls short for the LRK model for $\alpha\leq 1$, as the correlation decay pattern in the ground state is always algebraic.
@@
For this nonanalytic behavior of entanglement measure at $\mu_{f}=1$, we offer an alternate explanation in terms of the mode occupation probability $n_{k}$ for the soft mode ($k_{c}=\pi$) where the energy vanishes near criticality. Let us recall that $n_{k_{c}}=1/2$ would imply a contribution from multiple modes and therefore leads to higher entanglement entropy for the union of two subsystems, $S_{vN}^{A_{1}\cup A_{2}}$. As in $I_{A_{1}:A_{2}}$ in Eq. 9, $S_{vN}^{A_{1}\cup A_{2}}$ is being subtracted; this leads to a decrease in $I_{A_{1}:A_{2}}$. In contrast, $n_{k_{c}}=1,0$ will lead to a reduction of $S_{vN}^{A_{1}\cup A_{2}}$ and, consequently, leads to an increment of the mutual information.
@@
The occupation probability, $n_{k}=\langle\eta_{k}^{\dagger}\eta_{k}\rangle$ of the Bogoliubov modes, $\eta_{k}$, in the stationary state of the LRK model is obtained as
@@
where $\lambda^{i}_{\alpha}\left(k\right)(\lambda^{f}_{\alpha}\left(k\right))$ denotes the energy corresponding to the prequench (postquench) Hamiltonian. From Eq. 23, it can be shown that if we start with the noncritical initial state, i.e., $\mu_{i}=1.5$, in the limit $k\to\pi$, the occupation probability $n_{k}$ is given by
@@
as illustrated in Fig. 1(row 3). This explains the dips observed in $I_{A_{1},A_{2}}$ in the long-time stationary state for a noncritical-to-critical quench. In contrast, for the quench protocol involving the critical state, $n_{k}$ near the soft mode is
@@
This explains the peak observed in $I_{A_{1}:A_{2}}$ for the critical-to-critical quench protocol.
Going beyond the search for signatures of criticality in the bipartite information-theoretic measures such as mutual information and log-negativity, we now study the tripartite mutual information whose negative, zero, and positive values indicate perfectly delocalized or scrambled information, extensivity of mutual information, and redundancy of information [90; 91; 92]. For the completeness, the tripartite mutual information (TMI) is defined as
$I_{A_{1}:A_{2}:A_{3}}=I_{A_{1}:A_{2}}+I_{A_{1}:A_{3}}-I_{A_{1}:A_{2}\cup A_{3}},$ (26)
where $I_{A_{i}:A_{j}}$ is the mutual information between $A_{i}$ and $A_{j}$ [93]. It is known that for 1D noncritical systems, $I_{A_{1}:A_{2}:A_{3}}$ approaches zero in the large system limit [94]. For the same quench protocols used in the study of mutual information, TMI for the stationary state is presented in Fig. 2. Similar to the bipartite mutual information, the TMI for the stationary state obtained post critical-to-critical quench also exhibits a peak at $\mu=1$. In contrast, a nonanalytical behavior of TMI for the stationary state is visible for noncritical-to-critical quench. This indicates that the criticality associated with the stationary state can also be captured through tripartite mutual information, which is a measure of nonlocal correlations.
Figure 2: We plot the tripartite mutual information for different values of post quench $\mu$ by fixing the prequench $\mu_{i}$ at $\mu_{i}=1$ (left column) and $\mu_{i}=1.5$ (right column) for $\alpha=0$ (black squares), $\alpha=1$ (red circles) and $\alpha=2$ (blue triangles). The quench protocol for $\Delta$ is chosen to be the same as in Fig. 1
Moreover, a negative TMI for $\mu<1$ shows the presence of nonlocal information, but a redundancy of information for $\mu>1$ for $\alpha=1$. The short-range pairing term, i.e., $\alpha=2$, displays zero TMI everywhere except at the critical point, in agreement with the existing literature. An all-to-all pairing term, however, has an entirely different behavior where TMI is always positive, suggesting a redundancy of information, which peaks at the critical point, $\mu=1$.
Figure 3: Phase plot of effective central charge $c_{\rm eff}^{I}$ extracted from mutual information in the postquench $\Delta-\mu$ plane for both critical (left column) and noncritical (right column) initial state for $\alpha=0$ (top row), 1 (middle row) and 2 (bottom row). The parameter $\Delta_{i}$ is fixed at -1 for all the plots.
@@ IV.2 Signature of universality in stationary state
After establishing the nonanalytic behavior of mutual information (and log-negativity) for the postquench stationary state at the critical point of the postquench Hamiltonian, a natural question arises: Does the information of the effective central charge, $c_{\rm eff}^{I}$ ($c_{\rm eff}^{N}$), calculated from the scaling of $I_{A_{1}:A_{2}}$ ($\xi_{A_{1}:A_{2}}$), depend on the quench protocol? And is this effective central charge universal, i.e., independent of quench protocols in $\Delta$; if yes, is the central charge of the ground state conserved? To address these questions, we analyze both $c_{\rm eff}^{I}$ and $c_{\rm eff}^{N}$ for both of the quench protocols discussed in section IV.1. Here, it is useful to recall that central charge provides a powerful tool to understand the underlying CFT that describes the particular universality class [33]. For instance, the Ising universality class and the Luttinger liquid universality class are described by a completely different CFT and characterized by distinct central charges. The central charge gives a measure of the number of degrees of freedom or the “size” of the symmetry in the theory [95; 96].
Figure 3 displays $c^{I}_{\rm eff}$ as a function of the postquench parameters $\mu_{f}$ and $\Delta_{f}$. The columns of Fig. 3 represent $c^{I}_{\rm eff}$ for quenches starting with two distinct initial states, while the rows correspond to a varying range of $\alpha$. Let us focus on the first column, i.e., the quench from the critical state for different values of $\alpha$. For both $\alpha=0$ and $\alpha=2$, $c_{\rm eff}^{I}$ is independent of $\Delta$ quenches and retains the same value as the ground state for a critical-to-critical quench protocol. For other values of $\mu_{f}$, $c_{\rm eff}^{I}$ calculated from the stationary state takes the value of noncritical ground state, i.e., $0.5$ for $\alpha=0$ and 0 for $\alpha=2$  [^1] Note that $\Delta=0$ completely changes the Hamiltonian to a nearest neighbour model with no pairing term with effective central charge 0 for $\mu\neq 1$.. For $\alpha=1$, the effective central charge for the ground state at the critical point depends on the value of $\Delta$ and this character is retained in the postquench stationary state as well. Therefore, the universality can only be discussed in the cases of $\alpha=0$ and $\alpha=2$ from a critical-to-critical quench. It is clearly seen that $c_{\rm eff}^{I}$ calculated from the postquench stationary state conserves the central charge (see Table 1). This implies that the universality class for the critical ground state can also be inferred from the stationary state.
Table 2: Summary of effective central charge values calculated from postquench stationary state for critical-to-critical, noncritical-to-critical and critical-to-noncritical quenches of LRK model for different values of $\alpha$.
@@
Now for the second column of Fig. 3, the universal characteristic of the $c_{\rm eff}^{I}$ of the stationary state is reflected through the independence of $\Delta$ quenches for $\alpha=0$ and $\alpha=2$, while $\alpha=1$ remains $\Delta$ dependent. Note that even though the entanglement measures showed a non-analytic behavior at $\mu_{f}=1$ for the noncritical initial state, the effective central charge does not show any such transition for either $\alpha=0$ or 2. Therefore, despite the nonanalytic behavior of entanglement measures at $\mu_{f}=1$, we do not attribute it as a signature of criticality or a universal description in terms of the underlying CFT.
Therefore, to summarize, the signature of universality in terms of effective central charge is seen in the postquench stationary state for a critical-to-critical quench. The effective central charges values in this case are summarized in tabular form in Table 2, and are in agreement with the values in Table 1 calculated for the ground state.
Now, to investigate whether the long-range correlations in the stationary state are genuinely quantum in nature, we have evaluated the effective central charge $c_{\rm eff}^{N}$ extracted from the finite-size scaling of the log-negativity for the stationary state. While the details are in Appendix A, the findings of the effective central charge behavior and therefore universality remains the same as those calculated using mutual information.
@@ V Summary
In summary, we studied the quench dynamics of the long-range Kitaev chain to investigate the possible signatures of quantum criticality in the long-time stationary state for different values of the exponent $\alpha$ characterizing the range of $p$-wave pairing interaction. To probe the criticality, we consider the quantum information-based measures such as bipartite, and tripartite mutual information, and log-negativity. Our results show that a peak emerges at $\mu=1$ in bipartite mutual information, tripartite mutual information, and log-negativity for a critical-to-critical quench for all values of $\alpha$ studied in this work, exhibiting behavior similar to that of the critical ground state. This clearly indicates the presence of criticality in the stationary state. Importantly, in contrast to the short-range $XY$ model, the appearance of this peak cannot be attributed to the change in decay pattern of the fermionic correlation from exponential to algebraic at the critical point, as the fermionic correlators decay algebraically even at noncritical points for $\alpha\leq 1$. We argue that the peak in bipartite mutual information arises due to the change in the occupation probability of the Bogoliubov fermions in the soft mode at the critical point. To understand the longer range quantum information in a postquench stationary state, we studied the TMI, which not only captures the signature of criticality but also shows that quantum correlations are delocalized for the $\mu<1$ case and have redundancy for $\mu>1$ for $\alpha=1$. For pairing exponent $\alpha=2$, the mutual information is extensive for all values of $\mu$ except $\mu=1$, while for $\alpha=0$, the quantum correlations show redundancy.
We further show that for $\alpha=2$, long-range correlations develop only for the critical-to-critical quench. This result is obtained by analyzing the effective central charges extracted from bipartite mutual information and log-negativity scaling analysis. The effective central charge of the corresponding stationary state matches the same for the critical ground state, regardless of the quench in the strength of the $p$-wave pairing interaction, $\Delta$. This match indicates that the universality class for the critical ground state can also be inferred from the stationary state. In contrast, for $\alpha=0$, the effective central charge is nonzero for all quench protocols, leading to long-range correlations, except when the final Hamiltonian parameter $\Delta=0$. When $\Delta\neq 0$, the effective central charge is $1$ only for the critical-to-critical quench and is $1/2$ for all other quench protocols, consistent with the ground state. This consistency suggests that the stationary state for the critical-to-critical quench protocol belongs to the same universality class as the ground state.
For $\alpha=1$, the stationary state, similar to the ground state, cannot be described by any universality class, as the presence or absence of long-range correlations depends on the $\Delta$ quenches. However, long-range correlations develop for the critical-to-critical quench protocol, regardless of the $\Delta$ quenches.
@@ VI Acknowledgment
Appendix A Overview of central charge for postquench stationary state from log-negativity
Starting with the critical initial state, as shown in the top row, left column of Fig. 4, we find that for $\alpha=0$, $c_{\rm eff}^{N}=1/2$ for $\mu_{f}\neq 1$ and $\Delta_{f}\neq 0$, while $c_{\rm eff}^{N}=1$ for $\mu_{f}=1$ and $\Delta_{f}\neq 0$. In contrast, starting with a noncritical initial state, as shown in the top row, right column of Fig. 4, $c_{\rm eff}^{N}=1/2$ for all values of $\mu_{f}$ and $\Delta_{f}$, except for $\Delta_{f}=0$. This implies that for $\alpha=0$, $c_{\rm eff}^{N}=c_{\rm eff}^{I}$ for all quench protocols. For $\alpha=2$, similar to $c_{\rm eff}^{I}$, $c_{\rm eff}^{N}$ is nonzero and equal to $1/2$ only for a critical-to-critical quench protocol (with the exception of $\Delta_{f}=0$, where $c_{\rm eff}^{N}=0$). For all other quench protocols, $c_{\rm eff}^{N}=0$. This confirms that the long-range correlations associated with a nonzero effective central charge for both $\alpha=0$ and $\alpha=2$ in the stationary state are strictly quantum in nature. On the other hand, for $\alpha=1$, long-range correlations develop, irrespective of $\Delta$ quenches in the critical-to-critical quench protocol (see middle row, left column of Fig. 4). Interestingly, for this quench protocol, $c_{\rm eff}^{N}$ is slightly larger than $c_{\rm eff}^{I}$. In contrast, for the critical-to-noncritical quench protocol, the presence of long-range correlations in the stationary state depends on the $\Delta$ quenches. Specifically, if $\Delta_{f}\Delta_{i}\gtrsim 0$, both $c_{\rm eff}^{I}$ and $c_{\rm eff}^{N}$ are nonzero, whereas both are zero if $\Delta\Delta_{0}\lesssim 0$. For all the quench protocols where $\Delta_{f}\Delta_{i}\gtrsim 0$, $c_{\rm eff}^{N}>c_{\rm eff}^{I}$.
Figure 4: Phase plot of the effective central charge from log-negativity $c_{\rm eff}^{N}$ for both critical (first column) and noncritical (second column) initial state for $\alpha=0$ (first row), 1 (second row) and 2 (third row). The quench parameters are same as for Fig. 3.